40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 18 19:55:45 2009, rana, Configuration, Electronics, IP POS is back: ND filter gone, new resistors in IMG_0068.JPGIMG_0072.JPG
    Reply  Sun Oct 18 22:06:15 2009, rana, Configuration, Electronics, IP POS is back: ND filter gone, new resistors in Untitled.png
Message ID: 2112     Entry time: Sun Oct 18 22:06:15 2009     In reply to: 2110
Author: rana 
Type: Configuration 
Category: Electronics 
Subject: IP POS is back: ND filter gone, new resistors in 

I tried to compare the IP_POS time series with the IPANG and MC_TRANS but was foiled at first:

1) The IPANG scan rate was set to 0.5 second, so it doesn't resolve the pendulum motions well. Fixed in the .db file.

2) Someone had used a Windows/DOS editor to edit the .db file and it was filled with "^M" characters. I have removed them all using this command:   tr -d "\r" <ETMXaux.db > new.db

3) The MC_TRANS P/Y channels were on the MC Lock screen but had never been added to the DAQ. Remember, if there's a useful readback on an EPICS screen. its not necessarily in the frames unless you add it to the C0EDCU file. I have done that now and restarted the fb daqd. Channels now exist.

4) Changed the PREC of the IPPOS channels to 3 from 2.

5) changed the sign for the IBQPD (aka IPANG) so that bigger signal is positive on the EPICS screen.

Attachment 1: Untitled.png  41 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-